OpenStudy (spareb665):

Solve the equation 2^3x+1=5^-x and round your answer to two digits after the decimal sign.

OpenStudy (spareb665):



OpenStudy (stamp):

\[2^{3x+1}=5^{-x}\]\[(3x+1)ln2=-xln5\]\[3xln2+ln2=-xln5\]\[3xln2+xln5=-ln2\]\[x(3ln2+ln5)=-ln2\]\[x=-ln2/(3ln2+ln5)\]